# SLS-BRD: A system-level approach to seeking generalised feedback Nash equilibria

### v2

@@ 8            $k\coloneqq k+1$;
  • The convergence rate of the SLS-BRD is dominated by the slowest player: Whenever there exists a $p\in\mathcal{P}$ such that $L_{\widehat{BR}_{\Phi}}^{p}\gg L_{\widehat{BR}_{\Phi}}^{\tilde{p}}$ for all $\tilde{p}\in\mathcal{P}$, then $L_{\widehat{BR}_{\Phi}}\approx L_{\widehat{BR}_{\Phi}}^{p}$.
  • The norms $\|\bm{H}^{pp}\|_{2\to 2}=\|\bm{C}^{p}\bm{F_{\Phi}}^{p}+\bm{D}^{pp}\|_{2\to 2}^{2}$ increase with the FIR horizon $N$, specially for unstable games. In some cases, players might need to proportionally decrease the state-penalties $C^{p}$ to ensure (or improve) convergence.
@@
At the cost of interpretability, similar Lipschitz constants as in Eqs. (46)–(47) can be obtained also when $(\mathcal{X},\mathcal{U}_{\mathcal{G}},\mathcal{U}^{p})$ and $P$ are general, and when structural constraints are present. Moreover, we remark that Theorem 8 is obtained under the assumption that $\|\Phi_{x,N}\|_{F}^{2}<\gamma$ holds strictly. The best-response maps $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$ when this constraint is active, or when $\mathcal{W}$ is an ellipsoid, become quadratically-constrained quadratic programs and their Lipschitz properties are less intuitive.
IV Examples
@@
In the following, we demonstrate the SLS-BRD algorithm in two exemplary problems: Decentralised control of an unstable network and management of a competitive market. In both problems, we set the initial profile $\bm{K}_{0}=(\bm{\Phi_{u,0}}^{1},\ldots,\bm{\Phi_{u,0}}^{N_{P}})\bm{\Phi_{x,0}}^{-1}$ by projecting the zero-response $\bm{\hat{\Phi}_{u,0}}=\bm{0}$ into the feasible set. [^2] That is, $\bm{\Phi_{u,0}}=\operatorname*{arg\,min}_{\bm{\Phi_{u}}\in\bm{\mathcal{U}}_{\Phi}}\|\bm{\Phi_{u}}-\bm{\hat{\Phi}_{u,0}}\|_{\ell_{2}}^{2}$ given the feasible set $\bm{\mathcal{U}}_{\Phi}=\{\bm{\Phi_{u}}\in\ell_{2}[0,N):\bm{\Phi_{u}}^{p}\in U_{\Phi}^{p}(\bm{\Phi_{u}}^{-p}),\leavevmode\nobreak\ p\in\mathcal{P}\}$. A SLS-BRD routine is then simulated by having players update their policies through best-responses to the parametrisation of their rivals’ policies, assumed to be available. Due to numerical limitations, we interrupt the updates whenever the condition $\|\bm{\Phi}_{u,k}^{p}-\bm{\Phi}_{u,k-1}^{p}\|_{\ell_{2}}/\|\bm{\Phi}_{u,k}^{p}\|_{\ell_{2}}\leq 10^{-8}$ ($\forall p\in\mathcal{P}$) is satisfied.
IV-A Stabilisation of a bidirectional chain network
@@
$A=\begin{bmatrix}1&0.2&&\\ -0.2&\ddots&\ddots&\\ &\ddots&\ddots&0.2\\ &&-0.2&1\end{bmatrix};\leavevmode\nobreak\ \left\{B^{p}{=}\begin{bmatrix}\bm{0}_{6(p{-}1)\times 2}\\[6.45831pt] I_{2}\\[2.15277pt] \bm{0}_{6(3{-}p)\times 2}\end{bmatrix}\right\}_{p\in\mathcal{P}},$ (48)
where $A\in\mathbb{R}^{N_{x}\times N_{x}}$ and $B^{p}\in\mathbb{R}^{N_{x}\times N_{u}^{p}}$ with $N_{u}^{p}=2$ ($\forall p$). This game has unstable dynamics, since $\rho(A)=1.073>1$, however it is stabilisable for each $(A,B^{p})$. Moreover, the game is subjected to a noise process described by $w_{t}\sim\mathrm{Uniform}(\mathcal{W})$, $t\in\mathbb{N}$, defined over $\mathcal{W}=\{w_{t}\in\mathbb{R}^{N_{x}}:\|w_{t}\|_{\infty}\leq 1\}$. In this problem, players are interested in stabilising the game $\mathcal{G}_{\infty}^{\text{LQ}}$, while minimising their individual objective functionals,
$J^{p}(u^{p},u^{-p})=\mathrm{E}\left[\sum_{t=0}^{\infty}\Big{(}\|x_{t}\|_{2}^{2}+\beta^{p}\|u^{p}_{t}\|_{2}^{2}\Big{)}\right],$
[…]
@@
[…]
@@
enforcing $-10\leq[u_{t}^{p}]_{1}+[u_{t}^{p}]_{2}\leq 10$ for every $t\in\mathbb{N}$ and $p\in\mathcal{P}$. We assume that players design their state-feedback policies, $\bm{K}^{p}=\bm{\Phi_{u}}^{p}\bm{\Phi_{x}}^{-1}\in\bm{\mathcal{C}}^{p}$, considering a FIR horizon of $N=50$ and the constraints $(\Phi_{x,n}\in\mathcal{C}_{x,n})_{n=1}^{N}$ and $(\Phi_{u,n}^{p}\in\mathcal{C}_{u,n}^{p})_{n=1}^{N}$,
@@
$\mathcal{C}_{u,n}^{p}$ =$=\{\Phi_{u,n}^{p}\in\mathbb{R}^{N_{u}^{p}\times N_{x}}:\Phi_{u,n}^{p}=\texttt{Sp}({B^{p}}^{{\mkern-1.5mu\mathsf{T}}}A^{n-1})\odot\Phi_{u,n}^{p}\}.$
Under this setup, $\mathcal{G}_{\infty}^{\text{LQ}}$ belongs to the class of dynamic potential games (DPG, [45]) and the (unique) GFNE can be obtained in advance by solving a centralised optimisation problem.
In this experiment, we simulate a GFNE seeking procedure for each different value $\beta\in\{(10,40,10),(2,8,2),(0.4,1.6,0.4)\}$. The game $\mathcal{G}_{\infty}^{\text{LQ}}$ is executed alongside the updating of players’ policies according to some learning dynamics. The players seek $\varepsilon$-GFNE policies by adhering to the SLS-BRD routine (Algorithm 4) using their (approximately)best-response maps, $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$, with $\gamma=0.95$. The policies are updated simultaneously every $\Delta T=1$ stage with a learning rate of $\eta=1/2$.
The convergence of the SLS-BRD routine to the fixed-point $\bm{K}^{*}=\bm{\Phi_{u}}^{*}\bm{\Phi_{x}}^{-1}=(\bm{\Phi_{u}}^{1^{*}},\ldots,\bm{\Phi_{u}}^{N_{P}^{*}})\bm{\Phi_{x}}^{-1}$ is shown in Figure 3. The results demonstrate that the players’ policies are sufficiently close to the $\epsilon$-GFNE profile after $260$, $420$, and $510$ iterations for each respective weighting configuration. In each case, the (soft) FIR constraints are satisfied strictly after the initial profile (that is, $\|\Phi_{x,k,N}^{p}\|_{F}^{2}<0.95$, $k>2$, $p\in\mathcal{P}$). The Lipschitz properties of the global update rule $T_{\Phi}$ seem to follow the intuition provided in Section III-B: The convergence rate improves when players apply stronger penalties to their actions ($\beta=(10,40,10)$) when compared to that obtained by weaker control penalties ($\beta=(0.4,1.6,0.4)$). However, we stress that the Lipschitz constants from Theorem 8 do not consider structural constraints ($\bm{S}_{x}$ and $\bm{S}_{u}^{p}$, $\forall p\in\mathcal{P}$) and cannot be applied to this experiment. Regardless, the convergence rates are shown to be geometric, indicating that the best-response maps $\widehat{BR}_{\Phi}$ in these scenarios are contractive. If not interrupted, and disregarding numerical limitations, the SLS-BRD should continue to approach the fixed-point $\bm{K}^{*}$ at this rate.
@@
In Figure 4 (top), we show the relative distance between the individual updates $\{\bm{\Phi_{u,k}}^{p},\bm{\Phi_{u,k-1}}^{p}\}_{p\in\mathcal{P}}$ from each player. The updates are of similar magnitude for all players $p\in\mathcal{P}$, in each scenario, except for player $p=3$ which shows a slightly faster convergence. In general, these local changes become numerically negligible at a faster rate than the global convergence in Figure 3. Specifically, when the relative distance between updates approaches the aforementioned threshold of $10^{-8}$, the corresponding policy profile has become closer to the $\epsilon$-GFNE by a factor of $10^{-6}$. Finally, we consider the relative distances $\|\bm{\Phi_{x,k}}^{p}-\bm{\Phi_{x,k}}\|_{\ell_{2}}/\|\bm{\Phi_{x,k}}^{p}\|_{\ell_{2}}$, $k\in\mathbb{N}_{+}$, between the responses $\{\bm{\Phi_{x,k}}^{p}\}_{p\in\mathcal{P}}$ and $\bm{\Phi_{x,k}}=\bm{F_{\Phi}}\bm{\Phi_{u,k}}$ obtained from the system-level parametrisation associated with $\bm{\Phi_{u,k}}$ (Theorem 4). As shown in Figure 4 (bottom), these distances decrease at a similar rate and are relatively small since the initial stages of the game. The policies $\bm{K}_{k}=\big{(}\bm{\Phi_{u,k}}^{1}(\bm{\Phi_{x,k}}^{1})^{-1},\ldots,\bm{\Phi_{u,k}}^{N_{P}}(\bm{\Phi_{x,k}}^{N_{P}})^{-1}\big{)}$ thus approximate $(\bm{\Phi_{u,k}}^{1},\ldots,\bm{\Phi_{u,k}}^{N_{P}})\bm{\Phi_{x,k}}^{-1}$ as $k\to\infty$ and are expected to stabilise the system during this learning dynamics.
Figure 4: $N_{P}$-Chain game: Relative distance (in $\log$-scale) between updates $(\bm{\Phi_{u,k}}^{p},\bm{\Phi_{u,k+1}}^{p})$, top panels, and responses $(\bm{\Phi_{x,k}}^{p},\bm{\Phi_{x,k}})$, bottom panels.
The evolution of the game given the actions by each player is displayed in Figure 5 for the period $t\in(550,600]$, when the policy updates have been interrupted. The results demonstrate that the policies obtained by the SLS-BRD routine are capable of jointly stabilising the networked system, robustly against the random noise. These policies are also shown to satisfy the operational constraints: Each player’s actions become (roughly) symmetrical and satisfy $|[u_{t}]_{1}+[u_{t}]_{2}|\leq 5$, $t\in\mathbb{N}$, in all scenarios. The enforcement of this strategy highlights the conservativeness resulting from the robust operational constraints. Finally, note that the state- and control-trajectories from operating the system through these policies are similar for all $\beta$ configurations; however, the policies with $\beta=(2,8,2)$ and $\beta=(0.4,1.6,0.4)$ achieve a slightly better noise rejection at the expense of more aggressive control actions. In this experiment, the choice $\beta=(10,40,10)$ seems to be preferable as it converges faster while still achieving satisfactory performance.
IV-B Price competition in oligopolistic markets
@@
Define $\bm{x}=(\bm{x}^{1},\ldots,\bm{x}^{N_{P}})$, with $\bm{x}^{p}=(d^{p}(t)-\bar{d}_{base})_{t\in\mathbb{R}_{\geq 0}}$, and $\widetilde{B}^{p}=[\widetilde{B}^{p1}\leavevmode\nobreak\ \cdots\leavevmode\nobreak\ \widetilde{B}^{pN_{P}}]$, for all $p\in\mathcal{P}$. Considering a zero-order hold of inputs with period $\Delta t=1/4$ [days], the game $\mathcal{G}_{\infty}^{\text{LQ}}$ can be described by the discrete-time dynamics [^3] With a slight abuse of notation, we use $t$ for both the continuous-time ($x(t)$, $t\in\mathbb{R}_{\geq 0}$) and discrete-time ($x_{t}$, $t\in\mathbb{N}$) signals.
$x_{t+1}=Ax_{t}+\sum_{p\in\mathcal{P}}B^{p}u^{p}_{t}+w_{t}$
with $A=\exp(-\tau\Delta t)I_{N_{P}}$ and $\{B^{p}=-\frac{1}{\tau}(A-I_{N_{x}})\widetilde{B}^{p}\}_{p\in\mathcal{P}}$, and the noise process $\bm{w}=\big{(}-\frac{1}{\tau}(A-I_{N_{x}})v_{t}\big{)}_{t\in\mathbb{N}}$ [46]. The companies assume that the baseline demand fluctuations satisfy $w_{t}\in\mathcal{W}=\{w_{t}\in\mathbb{R}^{N_{x}}:\|w_{t}\|_{\infty}\leq 1\}$ for all $t\in\mathbb{N}$. Under this representation, each player’s objective is formulated as solving for a policy which minimises the functional
$J^{p}(u^{p},u^{-p})=\mathrm{E}\left[\sum_{t=0}^{\infty}\Big{(}\alpha^{p}\|x_{t}\|_{2}^{2}+\beta^{p}\|u^{p}_{t}\|_{2}^{2}\Big{)}\right],$
@@
$\mathcal{U}_{\mathcal{G}}$ =$=\big{\{}u_{t}\in\textstyle\prod_{p\in\mathcal{P}}\mathbb{R}^{N_{u}^{p}}:\textstyle\frac{1}{N_{P}\cdot\bar{u}_{\text{avg}}}\big{[}\mathbf{1}_{N_{u}}\leavevmode\nobreak\ {-}\mathbf{1}_{N_{u}}\big{]}^{{\mkern-1.5mu\mathsf{T}}}u_{t}\leq[1\leavevmode\nobreak\ 1]^{{\mkern-1.5mu\mathsf{T}}}\big{\}},$
@@
As in Section IV-A, we simulate an instance of the game $\mathcal{G}_{\infty}^{\text{LQ}}$ alongside the SLS-BRD routine (Algorithm 4) with players using (approximately)best-response maps, $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$, given $\gamma=0.95$. Policies are updated simultaneously every $\Delta T=1$ stage with learning rate $\eta=1/4$. Under the above setup, the game $\mathcal{G}_{\infty}^{\text{LQ}}$ is not a potential game and thus a GFNE cannot be easily computed in advance. Finally, we remark that solutions to this problem are neither unique nor necessarily admissible: Different initial profiles result in convergence to different fixed-points, which might be unfair to a subset of players.
The convergence of the SLS-BRD routine to a fixed-point $\bm{K}^{*}=\bm{\Phi_{u}}^{*}\bm{\Phi_{x}}^{-1}=(\bm{\Phi_{u}}^{1^{*}},\ldots,\bm{\Phi_{u}}^{N_{P}^{*}})\bm{\Phi_{x}}^{-1}$ is shown in Figure 6. Since the exact fixed-point to which the routine will converge is not known for this problem, we let $\bm{\Phi_{u}}^{*}\approx\bm{\Phi_{u,k_{f}}}$ for $k_{f}=828$ (when the policy updates are interrupted) and analyse the convergence with respect to this point. In this case, the iterates approach the fixed-point mostly at a geometric rate with the policies requiring roughly $800$ updates (or $200$ days in-game) before changes become numerically negligible. The rate of convergence is also observed to slow around $k=15$, when the generalised constraints ($\mathcal{U}_{\mathcal{G}}$) become active. We stress that the best-responses $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$ cannot be (globally) contractive in this case, as the set of fixed points $\Omega_{\mathcal{G}_{\infty}^{\Phi}}^{\varepsilon}$ is not a singleton.
@@
In Figure 7 (top), the relative distances between the individual updates $\{\bm{\Phi_{u,k}}^{p},\bm{\Phi_{u,k-1}}^{p}\}_{p\in\mathcal{P}}$ are displayed. As in the previous example, the updates are of similar magnitude for all players $p\in\mathcal{P}$ and they become numerically negligible at a faster rate than the global convergence in Figure 6. The convergence of the distance between updates is shown to slow considerably, especially for $p\in\{1,4\}$, around $k=50$. Thereafter, these relative distances continue to decrease at a steady rate. In Figure 7 (bottom), the relative distances $\|\bm{\Phi_{x,k}}^{p}-\bm{\Phi_{x,k}}\|_{\ell_{2}}/\|\bm{\Phi_{x,k}}^{p}\|_{\ell_{2}}$, $k\in\mathbb{N}_{+}$, between responses $\{\bm{\Phi_{x,k}}^{p}\}_{p\in\mathcal{P}}$ and $\bm{\Phi_{x,k}}=\bm{F_{\Phi}}\bm{\Phi_{u,k}}$ are shown. As before, these distances decrease at a similar rate and are relatively small since the initial stages of the game. The policies $\bm{K}_{k}=\big{(}\bm{\Phi_{u,k}}^{1}(\bm{\Phi_{x,k}}^{1})^{-1},\ldots,\bm{\Phi_{u,k}}^{N_{P}}(\bm{\Phi_{x,k}}^{N_{P}})^{-1}\big{)}$ thus approximate the stabilising profile $(\bm{\Phi_{u,k}}^{1},\ldots,\bm{\Phi_{u,k}}^{N_{P}})\bm{\Phi_{x,k}}^{-1}$ as $k\to\infty$ and are expected to be stabilising for all iterations.

### v3

@@ 8            $k\coloneqq k+1$;
  • The convergence rate of the SLS-BRD can be dominated by the slowest player: Whenever there exists a $p\in\mathcal{P}$ such that $L_{\widehat{BR}_{\Phi}}^{p}\gg L_{\widehat{BR}_{\Phi}}^{\tilde{p}}$ for all $\tilde{p}\in\mathcal{P}$, then $L_{\widehat{BR}_{\Phi}}\approx L_{\widehat{BR}_{\Phi}}^{p}$.
@@
At the cost of interpretability, similar Lipschitz constants as in Eq. (44) can also be obtained for general $(\mathcal{X},\mathcal{U}_{\mathcal{G}},\mathcal{U}^{p})$ and $P$, and when structural constraints are present. We stress that Theorem 4 is obtained under the assumption that $\|\Phi_{x,N}\|_{F}^{2}<\gamma$ holds strictly. The best-response maps $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$ when this constraint is active, or when $\mathcal{W}$ is an ellipsoid, are solutions to quadratically-constrained quadratic programs (QCQP) and their Lipschitz properties are less intuitive.
4 Examples
@@
In the following, we demonstrate the SLS-BRD algorithm in two exemplary problems: Decentralised control of an unstable network and management of a competitive market. In both problems, we set the initial profile $\bm{K}_{0}=(\bm{\Phi_{u,0}}^{1},\ldots,\bm{\Phi_{u,0}}^{N_{P}})\bm{\Phi_{x,0}}^{-1}$ by projecting the zero-response $\bm{\hat{\Phi}_{u,0}}=\bm{0}$ into the feasible set. A SLS-BRD routine is then simulated by having players update their policies through best-responses to the parametrisation of their rivals’ policies, assumed to be available. Due to numerical limitations, we interrupt the updates whenever the condition $\|\bm{\Phi}_{u,k}^{p}-\bm{\Phi}_{u,k-1}^{p}\|_{\ell_{2}}/\|\bm{\Phi}_{u,k}^{p}\|_{\ell_{2}}\leq 10^{-8}$ ($\forall p\in\mathcal{P}$) is satisfied.
4.1 Stabilisation of a bidirectional chain network
@@
$A=\begin{bmatrix}1&0.2&&\\ -0.2&\ddots&\ddots&\\ &\ddots&\ddots&0.2\\ &&-0.2&1\end{bmatrix};~{}\left\{B^{p}{=}\begin{bmatrix}\bm{0}_{6(p{-}1)\times 2}\\[6.45831pt] I_{2}\\[2.15277pt] \bm{0}_{6(3{-}p)\times 2}\end{bmatrix}\right\}_{p\in\mathcal{P}},$
where $A\in\mathbb{R}^{N_{x}\times N_{x}}$ and $B^{p}\in\mathbb{R}^{N_{x}\times N_{u}^{p}}$ with $N_{u}^{p}=2$ ($\forall p$). This game has unstable dynamics, since $\|A\|_{2}=1.073>1$, however it is stabilisable for each $(A,B^{p})$. Moreover, the game is subjected to a noise process described by $w_{t}\sim\mathrm{Uniform}(\mathcal{W})$, $t\in\mathbb{N}$, defined over $\mathcal{W}=\{w_{t}\in\mathbb{R}^{N_{x}}:\|w_{t}\|_{\infty}\leq 1\}$. In this problem, players are interested in stabilising the game $\mathcal{G}_{\infty}^{\text{LQ}}$, while minimising their individual objective functionals,
$J^{p}(\bm{u}^{p},\bm{u}^{-p})=\mathrm{E}\left[\sum_{t=0}^{\infty}\Big{(}\|x_{t}\|_{2}^{2}+\beta^{p}\|u^{p}_{t}\|_{2}^{2}\Big{)}\right],$
equivalent to Eq. (19) after setting $C^{p}=[I_{N_{x}}~{}\bm{0}_{N_{x}{\times}N_{u}}]^{{\mkern-1.5mu\mathsf{T}}}$, $D^{pp}=[\bm{0}_{N_{u}{\times}N_{x}}~{}\sqrt{\beta^{p}}I_{Nu}]^{{\mkern-1.5mu\mathsf{T}}}$, and $D^{p\tilde{p}}=0$ for all $\tilde{p}\in\mathcal{P}\backslash\{p\}$. The players’ actions are subjected to operational constraints, $\bm{u}^{p}\in U^{p}(\bm{u}^{-p})$, defined by the constraint sets
@@
$\mathcal{U}^{p}$ =$=\big{\{}u^{p}_{t}\in\mathbb{R}^{N_{u}^{p}}:\begin{bmatrix}\tfrac{1}{12}I_{N_{u}^{p}}\end{bmatrix}u^{p}_{t}\preceq\bm{1}_{N_{u}^{p}}\big{\}};$
@@
enforcing $\|u_{t}^{p}\|_{\infty}\leq 12$ for each $t\in\mathbb{N}$ and $p\in\mathcal{P}$ (Remark 2). We assume that players design their state-feedback policies, $\bm{K}^{p}=\bm{\Phi_{u}}^{p}\bm{\Phi_{x}}^{-1}\in\bm{\mathcal{C}}^{p}$, considering a FIR horizon of $N=50$ and the constraints $(\Phi_{x,n}\in\mathcal{C}_{x,n})_{n=1}^{N}$ and $(\Phi_{u,n}^{p}\in\mathcal{C}_{u,n}^{p})_{n=1}^{N}$,
@@
$\mathcal{C}_{u,n}^{p}$ =$=\{\Phi_{u,n}^{p}\in\mathbb{R}^{N_{u}^{p}\times N_{x}}:\Phi_{u,n}^{p}=\texttt{Sp}({B^{p}}^{{\mkern-1.5mu\mathsf{T}}}A^{n-1})\odot\Phi_{u,n}^{p}\},$
defined by setting the delay parameters $d_{a}=d_{c}=1$. Under this setup, $\mathcal{G}_{\infty}^{\text{LQ}}$ belongs to the class of dynamic potential games (DPG, [46]) and the (unique) GFNE can be obtained in advance by solving a centralised optimisation problem.
In this experiment, we simulate a GFNE seeking procedure for each value $\beta\in\{(10,40,10),(2,8,2),(0.4,1.6,0.4)\}$. The game $\mathcal{G}_{\infty}^{\text{LQ}}$ is executed alongside the updating of players’ policies according to some learning dynamics. The players are assumed to seek $\varepsilon$-GFNE policies by adhering to the SLS-BRD routine (Algorithm 4) using (approximately)best-response maps, $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$, with $\gamma=0.95$. The policies are updated simultaneously every $\Delta T=1$ stage with learning rate $\eta=1/2$.
The convergence of the SLS-BRD routine to the fixed-point $\bm{K}^{\star}=\bm{\Phi_{u}}^{\star}\bm{\Phi_{x}}^{-1}=(\bm{\Phi_{u}}^{1^{\star}},\ldots,\bm{\Phi_{u}}^{N_{P}^{\star}})\bm{\Phi_{x}}^{-1}$ is shown in Figure 3. The results demonstrate that the players’ policies are sufficiently close to the $\epsilon$-GFNE profile after $260$, $370$, and $425$ iterations for each respective weighting configuration. In each case, the (soft) FIR constraints are satisfied strictly after the initial profile (that is, $\|\Phi_{x,k,N}^{p}\|_{F}^{2}<0.95$, $k>2$, $\forall p\in\mathcal{P}$). Moreover, this fixed-point iteration seems to follow the behaviour discussed in Section 3.2: The convergence improves when players apply stronger penalties to their actions ($\beta=(10,40,10)$) when compared to that obtained by weaker control penalties ($\beta=(0.4,1.6,0.4)$). However, we stress that the Lipschitz constants from Theorem 4 do not consider structural constraints ($\bm{S}_{x}$, $\bm{S}_{u}^{p}$) and thus cannot be applied to this experiment. Regardless, the convergence is shown to be geometric, indicating that the best-response maps $\widehat{BR}_{\Phi}$ are contractive. If not interrupted, and disregarding numerical limitations, the SLS-BRD should continue to approach the fixed-point $\bm{K}^{\star}$ at this rate.
@@
Figure 4: $N_{P}$-chain game: Relative distance between the local updates $(\bm{\Phi_{u,k}}^{p},\bm{\Phi_{u,k-1}}^{p})$, top row, and responses $(\bm{\Phi_{x,k}}^{p},\bm{\Phi_{x,k}})$, bottom row.
@@
In Figure 4 (top), we show the relative distance between the individual updates $\{\bm{\Phi_{u,k}}^{p},\bm{\Phi_{u,k-1}}^{p}\}_{p\in\mathcal{P}}$ from each player. The updates are of similar magnitude for all players $p\in\mathcal{P}$, in each scenario, except for player $p=3$ which shows a slightly faster convergence. In general, these local changes become numerically negligible at a faster rate than the global convergence in Figure 3. Specifically, when the relative distance between updates approaches the aforementioned threshold of $10^{-8}$, the corresponding policy profile has become closer to the $\epsilon$-GFNE by a factor of $10^{-6}$. Finally, we consider the relative distances $\|\bm{\Phi_{x,k}}^{p}-\bm{\Phi_{x,k}}\|_{\ell_{2}}/\|\bm{\Phi_{x,k}}^{p}\|_{\ell_{2}}$, $k\in\mathbb{N}_{+}$, between the responses $\{\bm{\Phi_{x,k}}^{p}\}_{p\in\mathcal{P}}$ and $\bm{\Phi_{x,k}}=\bm{F_{\Phi}}\bm{\Phi_{u,k}}$ obtained from the system-level parametrisation associated with $\bm{\Phi_{u,k}}$ (Theorem 1). As shown in Figure 4 (bottom), these distances decrease at a similar rate and are relatively small since the initial stages of the game. The policies $\bm{K}_{k}=\big{(}\bm{\Phi_{u,k}}^{1}(\bm{\Phi_{x,k}}^{1})^{-1},\ldots,\bm{\Phi_{u,k}}^{N_{P}}(\bm{\Phi_{x,k}}^{N_{P}})^{-1}\big{)}$ thus approximate $(\bm{\Phi_{u,k}}^{1},\ldots,\bm{\Phi_{u,k}}^{N_{P}})\bm{\Phi_{x,k}}^{-1}$ as $k\to\infty$ and are thus expected to stabilise the system during this learning dynamics.
The evolution of the game given is displayed in Figure 5 for the period $t\in(550,600]$, after policy updates are interrupted. The results demonstrate that the policies obtained by the SLS-BRD routine are capable of jointly stabilising the networked system, robustly against the random noise. These policies are also shown to satisfy the operational constraints: Each player’s actions satisfy $\|u_{t}^{p}\|_{\infty}\leq 5$, $t\in\mathbb{N}$, in all scenarios. The enforcement of this strategy highlights the conservativeness resulting from the robust operational constraints. Finally, note that the state- and control-trajectories from operating the system through these policies are similar for all $\beta$ configurations; however, the policies with $\beta=(2,8,2)$ and $\beta=(0.4,1.6,0.4)$ achieve a slightly better noise rejection at the expense of more aggressive control actions. In this experiment, the choice $\beta=(10,40,10)$ seems to be preferable as it converges faster while still achieving satisfactory performance.
4.2 Price competition in oligopolistic markets
@@
Define $\bm{x}=(\bm{x}^{1},\ldots,\bm{x}^{N_{P}})$, with $\bm{x}^{p}=(d^{p}(t)-\bar{d}_{base})_{t\in\mathbb{R}_{\geq 0}}$, and $\widetilde{B}^{p}=[\widetilde{B}^{p1}~{}\cdots~{}\widetilde{B}^{pN_{P}}]$, for all $p\in\mathcal{P}$. Considering a zero-order hold of inputs with period $\Delta t=1/4$ [days], the game $\mathcal{G}_{\infty}^{\text{LQ}}$ can be described by the discrete-time dynamics [^1] With a slight abuse of notation, we use $t$ to index both the continuous-time ($x(t)$, $t\in\mathbb{R}_{\geq 0}$) and discrete-time ($x_{t}$, $t\in\mathbb{N}$) signals in this example.
$x_{t+1}=Ax_{t}+\textstyle\sum_{p\in\mathcal{P}}B^{p}u^{p}_{t}+w_{t}$
with $A=\exp(-\tau\Delta t)I_{N_{x}}$ and $\{B^{p}=-\frac{1}{\tau}(A-I_{N_{x}})\widetilde{B}^{p}\}_{p\in\mathcal{P}}$, and the noise process $\bm{w}=\big{(}-\frac{1}{\tau}(A-I_{N_{x}})v_{t}\big{)}_{t\in\mathbb{N}}$ [47]. The companies assume that the baseline demand fluctuations satisfy $w_{t}\in\mathcal{W}=\{w_{t}\in\mathbb{R}^{N_{x}}:\|w_{t}\|_{\infty}\leq 1\}$ for all $t\in\mathbb{N}$. Under this representation, each player’s objective is formulated as solving for a policy which minimises the functional
$J^{p}(\bm{u}^{p},\bm{u}^{-p})=\mathrm{E}\left[\sum_{t=0}^{\infty}\Big{(}\alpha^{p}\|x_{t}\|_{2}^{2}+\beta^{p}\|u^{p}_{t}\|_{2}^{2}\Big{)}\right],$
@@
$\mathcal{U}_{\mathcal{G}}$ =$=\big{\{}u_{t}\in\textstyle\prod_{p\in\mathcal{P}}\mathbb{R}^{N_{u}^{p}}:\begin{bmatrix}\tfrac{1}{N_{P}\cdot\bar{u}_{\text{avg}}}\bm{1}_{N_{u}}^{{\mkern-1.5mu\mathsf{T}}}\end{bmatrix}u_{t}\leq 1\big{\}}.$
@@
As in Section 4.1, we simulate an instance of the game $\mathcal{G}_{\infty}^{\text{LQ}}$ alongside the SLS-BRD routine (Algorithm 4) with players using (approximately)best-response maps, $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$, given $\gamma=0.95$. Policies are updated simultaneously every $\Delta T=1$ stage with learning rate $\eta=1/4$. Under the above setup, the game $\mathcal{G}_{\infty}^{\text{LQ}}$ is not a potential game and thus a GFNE cannot be easily computed in advance. Finally, we remark that solutions to this problem are not unique: Different initial profiles may result in convergence to different equilibrium profiles.